OpenStudy (anonymous):

What is the probability of tossing three coins and having exactly one of them show tails?

OpenStudy (anmolsingh):

the possible outcomes of such a toss are given as 2^(no of coins or tosses)=2^3=8 now there will be 3 times when each coin shows EXACTLY 1 TAIL. so the probability is 3/8

OpenStudy (amorfide):

anmol you are wrong... only one can show tails not all

OpenStudy (anonymous):

Thanks SOO much "AnmolSingh

OpenStudy (anmolsingh):

no amorfide I am right.... There willbe only 3 times when there will be 1 tails

OpenStudy (anmolsingh):

your welcome Mathew

OpenStudy (across):

When the sample space is small and you're lazy, you can always list all the possibilities and count: HHH HHT* HTH* HTT THH* THT TTH TTT 3/8

OpenStudy (anmolsingh):

well I was right:D @ amorfide

OpenStudy (amorfide):

my bad